OpenStudy (anonymous):

What is the simplified form of 8x^2/x^2+4x+4/4x/x^2-4

OpenStudy (anonymous):

it's a fraction on top of a fraction

OpenStudy (anonymous):

a/b/c=ac/b

OpenStudy (anonymous):

8x^2/x^2+4x+4/4x/x^2-4=8+4x+4x^2/4x-4=8+4x+x-4=5x+4

OpenStudy (anonymous):

waht do u think? am i clear